Aluminum bromate is a hydrated compound (that is, there are water molecules associate with the aluminum bromate). If you heat 15.0801 g of this salt, it loses its water and the mass remaining is found to be 11.1626 g. What is the number of water molecules in the molecular formula of the hydrate and what is the molecular formula of aluminum bromate hydrate?
1
Expert's answer
2012-10-08T10:24:47-0400
The molecular mass of aluminium bromate is 410.7 g/mol. The one mol of
aluminium bromate hydrate can produce x moles of water:
Al(BrO3)3*xH2O =
Al(BrO3)3 + x H2O
The mass of the water is 15.0801 - 11.1626 = 3.9175
g
The amount of aluminium bromate is 11.1626/410.7 = 0.0272 mol
The amount
of water is 3.9175/18 = 0.2176 mol
The number of water molecules is
x =
n(H2O)/n(Al(BrO3)3) = 0.2176 / 0.0272 = 8
The molecular formula of aluminium
bromate hydrate is
Al(BrO3)3*8H2O